A mathematics teacher buys some books for her A-Level and IGCSE students. A-Level books cost $10 each and IGCSE books $15 each. She spends a total of $1800, buying a total of 160 books in all.
1) Write two simultaneous equations to express this information, defining your symbols clearly.
2) Solve your equations to find out how many of each type of book she buys.

Answer by Boreal(15235)

A-level is x
IGCSE is y
x+y=160
10x+15y=1800
-10x-10y=-1600, multiplying the top by -10
5y=200
y=40 books( IGCSE) that are $15 each or $600
x=120 books (A-level) that are $10 each or $1200